OpenStudy (anonymous):

change each fraction to decimals 4/5 =?

OpenStudy (anonymous):

4/5 = 0.8

OpenStudy (anonymous):

To make it into decimal, the base should be 10,100..... powers of 10. 4*2/5*2 = 8/10=0.8